OpenStudy (anonymous):

Use the slope and y-intercept to graph each equation. y+2=5x-4 What is the slope and y-intercept, and how do i find them?

jimthompson5910 (jim_thompson5910):

First solve for y y+2=5x-4 y=5x-4-2 y = 5x - 6

jimthompson5910 (jim_thompson5910):

The equation y = 5x - 6 is in the form y = mx+b where m = 5 (slope) b = -6 (y-intercept)

jimthompson5910 (jim_thompson5910):

to graph this, you plot the point (0, -6), which is the y-intercept as a point then you go up 5 units and over to the right 1 unit to get to the next point (1, -1)

jimthompson5910 (jim_thompson5910):

you now have the two points: (0, -6) and (1,-1) plot them both and draw a straight line through them to graph y = 5x - 6

OpenStudy (anonymous):

Thank you jim :)

jimthompson5910 (jim_thompson5910):

you're welcome